OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 -1 1 31

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (just_one_last_goodbye):

Step: Solve6x+y=4for y: 6x+y+−6x=4+−6x(Add -6x to both sides) y=−6x+4 Step: Substitute−6x+4foryin5x+y−7=0: 5x+y−7=0 5x+−6x+4−7=0 −x−3=0(Simplify both sides of the equation) −x−3+3=0+3(Add 3 to both sides) −x=3 −x −1 = 3 −1 (Divide both sides by -1) x=−3 Step: Substitute−3forxiny=−6x+4: y=−6x+4 y=(−6)(−3)+4 y=22(Simplify both sides of the equation) Answer: y=22 and x=−3

OpenStudy (anonymous):

so -1?

OpenStudy (trojanpoem):

Is this cremer's rule ?

OpenStudy (anonymous):

its solving linear systems algebraically

OpenStudy (anonymous):

thats all i know...

OpenStudy (trojanpoem):